Find the P-value for the indicated hypothesis test.
A manufacturer claims that fewer than 6% of its fax machines are defective. In a random sample of 97 such fax machines, 5% are defective. Find the P-value for a test of the manufacturer's claim.

Answer by stanbon(57290)

Find the P-value for the indicated hypothesis test.
A manufacturer claims that fewer than 6% of its fax machines are defective.
This implies to have a left-tail test.
--------------------------------------
In a random sample of 97 such fax machines, 5% are defective.
Find the z-value of that sample proportion:
z(0.05) = (0.05-0.06)/sqrt[0.05*0.95/97] = -0.4520
------------------------
Find the P-value for a test of the manufacturer's claim.
p-value = P(z < -0.4520) = tcdf(-100,-0.4520 when df = 96) = 0.3262
============
I used a TI-84 to get the p-value
============
Cheers,
Stan H.